### v1

@@ Abstract
We formalize what it means to have conceptual knowledge about a statistical decision-making environment. Such knowledge tells agents about the structural relationships among unknown, payoff-relevant states. It allows agents to represent states as combinations of features. Conceptual knowledge is more valuable when states are more “reducible”: when their prior variances are explained by fewer features. Its value is non-monotone in the quantity and quality of available data, and vanishes with infinite data. Agents with deeper knowledge can attain the same welfare with less data. This is especially true when states are highly reducible.
@@ 1 Introduction
Humans use “concepts”—mental representations—to make sense of the world and take actions in it (Murphy,, 2002). Concepts help us describe the features of, and relationships among, objects in our environment. In this paper, we ask: when and why is conceptual knowledge “valuable?”
For example, suppose a farmer wants to learn the effects of applying fertilizers to his crops. He does so by trialing a combination of fertilizers and observing the resulting yield. The farmer knows the fertilizers supply nitrogen, a nutrient that helps plants grow. This allows him to mentally represent each fertilizer’s effect as the sum of a common “nitrogen effect” and an idiosyncratic effect. The farmer also knows the nitrogen effect explains more of fertilizers’ overall effects than do the idiosyncratic effects. So, in his trial, he combines fertilizers so as to isolate the nitrogen effect and hold constant the idiosyncratic effects. This makes the trial maximally informative: it leads to the largest reduction in his uncertainty about fertilizers’ overall effects.
If the farmer did not know about nitrogen (or, more generally, the concept of a nutrient), then he would not know to isolate the effect of supplying it. Instead he would view fertilizers as “black boxes,” knowing that they help plants grow but not knowing why. So he would have to learn about each fertilizer separately, making his trials less informative. His conceptual knowledge—his ability to represent fertilizers’ effects as sums of common and idiosyncratic effects—empowers him to run more informative trials. This information is valuable because it helps him make better decisions about which fertilizers to use. In contrast, his conceptual knowledge is valuable because it helps him collect more valuable information. This notion of “value” is what we study.
Quantifying the value of conceptual knowledge is important for designing interventions that give people information. It can be beneficial to give them knowledge that supplements their information and helps them interpret it (Sankar et al.,, 2025). If we know the value of such knowledge, then we can design better interventions that improve people’s lives more cost-effectively.
@@
Understanding how conceptual knowledge affects inferences is also important for understanding humans’ role in the age of artificial intelligence (AI). Humans (currently) have an edge in using concepts to unite seemingly unrelated phenomena: Newton isolated a concept (gravity) that unites apples falling on earth with the orbits of other planets; Watson and Crick isolated a concept (DNA) that unites crime scene investigation with the origins of domesticated rice; Bernoulli isolated a concept (risk aversion) that unites choices in poker with choices between insurance policies. These and other concepts allow humans to learn from limited data. In contrast, AI relies on recognizing patterns in large, rich sets of data. By studying what conceptual knowledge is, and when and why it is valuable, we can better allocate inferential tasks between humans and AI.
@@
This paper makes three major contributions. First, we formalize what it means to have conceptual knowledge about a statistical decision-making environment. Such knowledge refers to understanding the structural relationships among unknown, payoff-relevant states. It is distinct from the information one gains about these states from collecting data; it improves one’s ability to interpret data and makes them more informative. This distinction has been recognized by others (e.g., Marinacci,, 2015), but has not been formalized in a tractable and transparent way.
Second, we use our formalization of conceptual knowledge to define and characterize its value. Our definition builds on that for the instrumental value of information (Howard,, 1966; Raiffa and Schlaifer,, 1961): whereas information is valuable because it leads to better decisions, knowledge is valuable because it leads to better information. This definition is novel. It leads to a precise yet intuitive measure of the value of conceptual knowledge: the welfare gain from knowing how states relate structurally and using that knowledge to collect data optimally.
Third, we define the “units” of conceptual knowledge as the depth of one’s understanding of structural relationships. This definition is also novel. It allows us to compare the marginal values of “deepening” one’s knowledge and having more data. It also identifies conceptual knowledge as an economic good one may acquire in the same way data are goods one acquires. In this way, we advance the literature on learning and information acquisition that treats structural knowledge as fixed and minimally restrictive (e.g., Bardhi,, 2024; Callander,, 2011; Schwartzstein,, 2014).
This paper also contributes to the literatures on model-based learning, model uncertainty and mis-specification, and human and machine reasoning. We explain these contributions later.
@@
Section 3 extends the example to a more general setting. We consider a Bayesian agent who makes a statistical decision. His environment contains a collection of unknown, real-valued states. He learns about these states from noisy data. Then he takes real-valued actions. His loss equals the mean squared difference between the actions and states. The agent takes the actions that minimize his expected posterior loss.
@@
The agent’s prior on the states encodes his conceptual knowledge. This knowledge tells him how the states relate structurally. It allows him to represent them as linear combinations of “features.” These features form a basis for the Euclidean space of state vectors. They are eigenvectors of the state vector’s prior variance matrix. The corresponding eigenvalues capture features’ influences on the states: an eigenvalue is larger when the corresponding feature contributes more to the states’ prior variances.
@@
We say states are more “reducible” when they are influenced by fewer features. This happens when the eigenvalues of the prior variance matrix are more spread out. For example, if one eigenvalue is much larger than the others, then the state vector is likely to belong to a one-dimensional subspace of the many-dimensional state space. The more spread out are the eigenvalues, the more the agent can “reduce” states by representing them as low-dimensional combinations of high-dimensional features. This dimension reduction is what makes conceptual knowledge valuable.
@@
Section 4 contains definitions and preliminary results that we draw upon later. For example, we define the “value” of the agent’s data, derive sharp bounds on this value (see Proposition 1), and define what it means for eigenvalues to be “more spread out.”
We define the value of conceptual knowledge in Section 5. First, we suppose the agent collects an “optimal sample” that leads to the least posterior expected loss among all samples of a given size. This sample contains information about the most influential features only (see Proposition 2). Its design relies on the agent’s knowledge of how the states relate structurally.
Next, we consider a counterfactual agent who does not know how the states relate structurally. This “naïve” agent has no reason to believe the states covary or have different prior variances. So he assumes they do not. Consequently, when he collects an optimal sample, he tries to learn about every state equally. This is because he cannot identify features with different influences on the states, precluding him from focusing on the most influential features.
We compute the difference between the values of the true and naïve agent’s optimal samples, and we call this difference the “value of conceptual knowledge.” It captures the welfare gain from knowing how the states relate structurally and using that knowledge to collect data optimally.
@@
Our first main result (Theorem 1) says that conceptual knowledge is more valuable when states are more reducible. If a few features contribute most of the states’ prior variances, then the agent gains a lot from identifying those features and collecting data on them (i.e., “asking the right questions”). In contrast, if every feature contributes equally, then he gains nothing from identifying those features because he collects the same data that he would if he was naïve.
@@
Our second main result (Theorem 2) says that the value of conceptual knowledge (i) is non-monotone in the agent’s access to data and (ii) vanishes when he has infinite data. If the agent has more data, then he can learn more about the features on which he collects data, raising the gain from identifying the most influential features. However, he can also collect data on more features, lowering the gain from identifying the most influential features. The first effect dominates the second precisely when the sample is sufficiently small. As it becomes arbitrarily large, the agent’s posterior becomes independent of his prior, and so the conceptual knowledge embedded in his prior becomes irrelevant and loses its instrumental value.
@@
In Section 6, we extend our measure of the value of conceptual knowledge to one of “deeper” knowledge. We suppose the agent can identify some, but not all, of the most influential features, and refer to the “depth” of his knowledge as the number he can identify. Our third main result (Theorem 3) says that deeper knowledge is weakly more valuable. However, if the agent’s knowledge is sufficiently deep, then deepening it further yields no additional value because it does not change the data he collects.
@@
Finally, in Section 7, we study the trade-off between having deeper knowledge or more data. We measure the agent’s welfare by the least posterior expected loss he can attain given the depth of his knowledge and his access to data. Our fourth main result (Theorem 4) says that if he has deeper knowledge or the states are more reducible, then he can attain the same welfare with less data. This is because he can design better samples and extract more value from each observation, lowering the number he needs to attain a given welfare target.
@@
Section 8 discusses our modeling assumptions and related literature. Section 9 concludes. Appendix A contains additional discussions and results. Appendix B contains proofs of our mathematical claims.
@@ 2 An illustrative example
of using $w_{1}\in\mathbb{R}$ more units of fertilizer 1 and $w_{2}\in\mathbb{R}$ more units of fertilizer 2.111 We interpret negative values of $w_{k}$ as using less of fertilizer $k$ than the farmer uses currently. The vector $w=(w_{1},w_{2})$ has Euclidean length $\lVert w\rVert=1$ and the error $u\in\mathbb{R}$ is independently normally distributed with variance $\sigma_{u}^{2}>0$. It captures the randomness in $y$ due to variation in unobserved factors.
@@
The farmer knows the two fertilizers supply equal amounts of nitrogen, a nutrient that helps plants grow. This knowledge is purely conceptual: it comes from recognizing that fertilizers supply nutrients and from understanding the mechanisms through which plants grow, rather than from observing data on plants’ growth. It allows him to mentally represent each effect $\theta_{k}$ as the sum of a common “nitrogen effect” and an idiosyncratic effect. He encodes these effects by the scalars
@@
These vectors form an orthonormal basis for the Euclidean space $\mathbb{R}^{2}$ containing $\theta$. The common and idiosyncratic effects $\gamma_{1}$ and $\gamma_{2}$ are the coordinates of $\theta$ over this basis. The farmer does not know these coordinates, but he knows $\gamma_{1}$ contributes more to the prior variances of $\theta_{1}$ and $\theta_{2}$ than does $\gamma_{2}$. So he assumes $\gamma_{1}$ and $\gamma_{2}$ are independently distributed with variances $\lambda_{1}=\sigma^{2}(1+\rho)$ and $\lambda_{2}=\sigma^{2}(1-\rho)$. The sum
@@
of this sum contributed by $\gamma_{1}$. This share equals $1/2$ when $\rho=0$, in which case $\gamma_{1}$ and $\gamma_{2}$ contribute equally. It equals one in the limit as $\rho\to 1$, in which case only $\gamma_{1}$ contributes. The larger is $\rho$, the more likely is $\theta$ to belong to the one-dimensional subspace of $\mathbb{R}^{2}$ spanned by $v_{1}$.
@@
The farmer’s data ${\mathcal{S}}\equiv\{(w,y)\}$ comprise the vector $w=(w_{1},w_{2})$ and outcome $y$. These data are valuable insofar as they make the farmer’s beliefs about $\theta$ more precise. We measure the value of $\mathcal{S}$} via the mean difference
between the prior and posterior variances of $\theta_{1}$ and $\theta_{2}$. This difference is largest when $w=\pm v_{1}$ and is smallest when $w=\pm v_{2}$ (see Proposition A2). For example, choosing $w=v_{1}$ makes $y=\gamma_{1}+u$ a “pure signal” of $\gamma_{1}$. This makes $\mathcal{S}$} maximally valuable because it provides information about the component of $\theta$ with the most prior variance, leading to the largest difference between prior and posterior variances. In contrast, choosing $w=v_{2}$ makes $y=\gamma_{2}+u$ a pure signal of $\gamma_{2}$. This makes $\mathcal{S}$} minimally valuable because it provides information about the component of $\theta$ with the least prior variance, leading to the smallest difference between prior and posterior variances.222 In general, the data $\mathcal{S}$} are most valuable when they contain information about the components of $\theta$ with the most prior variance. We formalize and prove this claim in Sections 4.3 and 4.4.
@@
The farmer attains $\pi^{*}$ by choosing $w=\pm v_{1}$. Doing so relies on his conceptual knowledge about fertlizers’ effects: his ability to identify the common and idiosyncratic components of each effect $\theta_{k}$. If the farmer lacked this ability, then he would have no reason to believe $\theta_{1}$ and $\theta_{2}$ were correlated. So he would assume $\rho=0$ and the data would have maximal value
@@
between $\pi^{*}$ and $\pi^{(0)}$ captures the value of the farmer’s conceptual knowledge: the value of knowing how $\theta_{1}$ and $\theta_{2}$ relate structurally, and using this knowledge to collect data optimally.
The value $\Pi$ of the farmer’s conceptual knowledge is larger when $\rho$ is larger.333 We have $\pi^{*}=\frac{(1+\rho)^{2}\sigma^{4}}{2\mathopen{}\mathclose{{\left((1+\rho)\sigma^{2}+\sigma_{u}^{2}}}\right)}\quad\text{and}\quad\pi^{(0)}=\frac{\sigma^{4}}{2(\sigma^{2}+\sigma_{u}^{2})}$ by Proposition A2 and the definition of $\pi^{(0)}$. So $\partial\pi^{*}/\partial\rho>0$ and $\partial\pi^{(0)}/\partial\rho=0$, from which it follows that $\partial\Pi/\partial\rho>0$. ,444 For example, the correlation $\rho$ will be close to one when the fertilizers supply nitrogen only, and close to zero when their nutrient profiles are very different. Intuitively, the more likely it is $\theta$ that belongs to a low-dimensional subspace, the more precision can be gained by identifying that subspace and focusing on it when collecting data. We formalize this intuition in Section 5, and generalize it to a setting in which $\mathcal{S}$} has arbitrary size and $\theta$ has arbitrary length. In this setting, conceptual knowledge is more valuable when the eigenvalues of the prior variance matrix $\operatorname{\mathbb{V}}(\theta)$ are more spread out (see Theorem 1). Hence raising $\rho$ raises $\Pi$: it raises $\lambda_{1}=(1+\rho)\sigma^{2}$ and lowers $\lambda_{2}=(1-\rho)\sigma^{2}$ without changing their mean $(\lambda_{1}+\lambda_{2})/2=\sigma^{2}$.

### v2

@@ Abstract
We study the instrumental value of conceptual knowledge when making statistical decisions. Such knowledge tells agents how unknown, payoff-relevant states relate. It is distinct from the statistical knowledge gained from observing signals of those states. We formalize this distinction in a tractable framework used by economists and statisticians. Conceptual knowledge is valuable because it empowers agents to design more informative signals. It is more valuable when states are more “reducible”: when they can be explained with fewer common concepts. Its value is non-monotone in the number of signals and vanishes when agents have infinitely many signals. Agents who know more concepts can attain the same payoffs with fewer signals. This is especially true when states are highly reducible.
@@ 1 Introduction
Humans use mental models to make sense of the world (Johnson-Laird,, 1983). The building blocks of these models are “concepts”: mental representations we use to describe objects in our environment and how they relate (Murphy,, 2002). Understanding these relationships empowers us to use information about one object to draw inferences about another (Mitchell,, 2021).
This paper studies the interaction between (i) the conceptual knowledge embedded in mental models and (ii) the statistical knowledge inferred from data. We ask: when and why is conceptual knowledge instrumentally “valuable”?
For example, suppose a farmer wants to learn which fertilizers to apply to his crops. He views fertilizers as “black boxes,” knowing that they help crops grow but not why. He does not know how fertilizers’ effects relate and cannot extrapolate one from another’s. So when he tries different fertilizers to learn their effects, he has to try them separately.
Now suppose the farmer knows fertilizers supply nitrogen, a nutrient that helps crops grow. This tells him how fertilizers’ effects relate: they share a common “nitrogen component.” He can use this component to extrapolate one fertilizer’s effect from another’s. Moreover, when he tries different fertilizers to learn their effects, he can combine them to isolate the nitrogen component. This is better than trying each fertilizer separately because it allows him to learn their effects from one trial rather than many.
Nitrogen is a mental construct—the farmer cannot see it. All he can see are the effects of trying different fertilizers. But his conceptual knowledge of nitrogen allows him to learn more efficiently. It empowers him to run trials that are more informative and instrumentally valuable. How much more valuable is a quantity we define and characterize in this paper. We call this quantity the “value of conceptual knowledge.”
Quantifying this value is important for designing interventions that give people information. In empirical work motivating and guided by this paper (Sankar et al.,, 2025), we experimentally study the effect of teaching concepts to farmers. We compare farmers given concepts and data to farmers given data only, and find that farmers given concepts make more profitable decisions. These results provide empirical support for the theoretical predictions made in this paper. The results suggest a way to improve interventions that give people data: we should also give them conceptual tools to interpret data. By explaining how these tools work and quantifying their value, this paper helps in designing interventions that improve people’s lives more cost-effectively.
@@
Quantifying the value of conceptual knowledge is also important for comparing human and artificial intelligence (AI). Humans currently have an edge in using concepts to unite seemingly unrelated phenomena: Newton isolated a concept (gravity) that unites apples falling on Earth with the orbits of other planets; Watson and Crick isolated a concept (DNA) that unites crime scene investigation with the origins of domesticated rice; Bernoulli isolated a concept (risk aversion) that unites choices in poker with choices between insurance policies. These and other concepts allow humans to learn from limited data (Tenenbaum et al.,, 2011). In contrast, AI relies on recognizing patterns in large, rich sets of data (Goodfellow et al.,, 2016; Halevy et al.,, 2009). By understanding what conceptual knowledge is, and when and why it is valuable, we can better allocate inferential tasks between humans and AI.
@@
This paper makes three major contributions. First, we distinguish two types of knowledge: statistical and conceptual. Statistical knowledge comes from observing patterns in data; conceptual knowledge tells us how data are generated and what patterns to expect ex ante. We formalize this distinction in a tractable framework used by economists and statisticians. This framework gives us a transparent model of how statistical and conceptual knowledge interact.
Second, we use our framework to define and characterize the value of conceptual knowledge. Our definition builds on that of the instrumental value of information (Howard,, 1966; Raiffa and Schlaifer,, 1961): whereas information is valuable because it leads to better decisions, conceptual knowledge is valuable because it leads to better information. We quantify how much better and study this quantity’s comparative statics. Our framework admits a closed-form expression for the value of optimally acquired information, allowing us to analyze and clarify how and why this value changes. Moreover, our focus on a specific decision problem allows us to generate insights and intuitions that are not immediate from studying abstract problems (as in, e.g., Blackwell, (1951, 1953) and Whitmeyer, (2025)). [^1] For example, our non-monotonicity result (Theorem 2) could not be derived from Blackwell, (1951, 1953) or Whitmeyer,’s (2025) frameworks without imposing more structure on them.
Third, we use our framework to formalize what it means to have “deeper” conceptual knowledge. This allows us to compare the marginal values of having deeper knowledge or more data. It also identifies conceptual knowledge as an economic good one may acquire in the same way data are goods one acquires. In this way, we advance the literature on learning and information acquisition that treats conceptual knowledge as fixed and minimally restrictive (e.g., Bardhi,, 2024; Callander,, 2011; Schwartzstein,, 2014).
This paper also connects to the literatures on the value of information, model-based inference, human cognition, and human-machine comparisons. We discuss these literatures and our contributions to them after presenting our main results.
@@
Section 3 extends the example to a more general setting. We consider a Bayesian agent who makes a statistical decision. His environment contains a collection of unknown, real-valued states. He learns about these states from noisy signals. Then he takes real-valued actions. His loss equals the mean squared difference between the actions and states. He takes the actions that minimize his posterior expected loss.
@@
The signals give the agent statistical knowledge. His prior on the states encodes his conceptual knowledge. This knowledge tells him how states relate. It allows him to mentally represent states as combinations of “concepts,” which we model as eigenvectors of the state vector’s prior variance matrix. The corresponding eigenvalues index concepts’ explanatory power: an eigenvalue is larger when the corresponding concept explains more of the states’ prior variances.
@@
We say states are more “reducible” when their prior variances are explained by fewer common concepts. This happens when the eigenvalues of the prior variance matrix are more spread out. For example, if one eigenvalue is much larger than the others, then the state vector is likely to be close to a one-dimensional subspace of the many-dimensional state space. The more spread out are the eigenvalues, the more the agent can “reduce” the state vector by representing it as a low-dimensional combination of high-dimensional concepts. This dimension reduction is what makes conceptual knowledge valuable.
@@
Section 4 contains preliminary results that we draw upon in later sections. We characterize the instrumental “value” of the agent’s signals, derive sharp bounds on this value (see Proposition 1), and formalize what it means for eigenvalues to be “more spread out.”
We define and characterize the “value of conceptual knowledge” in Section 5. First, we suppose the agent designs an “optimal sample” containing the most valuable signals. This sample focuses on the concepts with the most explanatory power (see Proposition 2). Next, we consider a counterfactual “naïve” agent who does not know which concepts have more or less explanatory power. This prevents him from focusing on the concepts with the most power. As a result, his optimal sample is less valuable than the conceptually knowledgeable agent’s. We derive the values of optimal samples with and without conceptual knowledge, and call their difference the “value” of such knowledge. This difference equals the payoff gain from having conceptual knowledge and using it to design a more valuable sample.
@@
Our first main result (Theorem 1) says that conceptual knowledge is more valuable when states are more reducible. If the states can be explained by a few common concepts, then the agent gains a lot from identifying those concepts and focusing on them when he designs signals (i.e., “asking the right questions”). In contrast, if every concept has the same explanatory power, then the agent gains nothing from identifying those concepts because he designs the same sample that he would if he was naïve.
@@
Our second main result (Theorem 2) says that the value of conceptual knowledge (i) is non-monotone in the number of signals and (ii) vanishes when there are infinitely many signals. If the agent can observe more signals, then he can learn more about the concepts he focuses on, raising the payoff gain from knowing which to focus on. However, having more signals also prompts him to broaden his focus, lowering the gain from knowing which concepts to focus on. The first effect dominates the second when the number of signals is sufficiently small. As it becomes arbitrarily large, the agent’s posterior becomes independent of his prior, and so the conceptual knowledge embedded in his prior becomes irrelevant and loses its instrumental value.
@@
In Section 6, we extend our measure of the value of conceptual knowledge to one of “deeper” knowledge. We suppose the agent knows some, but not all, of the relevant concepts, and refer to the “depth” of his knowledge as the number he knows. Our third main result (Theorem 3) says that deeper conceptual knowledge is weakly more valuable. However, if the agent knows enough concepts, then knowing more yields no additional value because it does not change the optimal sample he designs.
@@
Finally, in Section 7, we study the trade-off between conceptual and statistical knowledge. Our framework gives us a precise language for describing this trade-off: is the agent better off knowing more concepts or having more signals? Our fourth main result (Theorem 4) says that if the agent knows more concepts, then he can attain the same welfare with fewer signals, especially when states are highly reducible. This is because he can design better samples and extract more value from each signal, lowering the number he needs to attain a given welfare target.
@@
Section 8 discusses related literature. Section 9 concludes. Appendix A contains additional discussions and results. Appendix B contains proofs of our mathematical claims.
@@ 2 An illustrative example
of using $w_{1}\in\mathbb{R}$ more units of fertilizer 1 and $w_{2}\in\mathbb{R}$ more units of fertilizer 2.222We interpret negative values of $w_{k}$ as using less of fertilizer $k$ than the farmer uses currently. The vector $w=(w_{1},w_{2})$ has Euclidean length $\lVert w\rVert=1$ and the error $u\in\mathbb{R}$ is independently normally distributed with variance $\sigma_{u}^{2}>0$. [^3] We normalize $\lVert w\rVert=1$ so that only the direction of $w$ (and not its magnitude) affects the informativeness of $y$. It captures the randomness in $y$ due to variation in unobserved factors.
The farmer’s data ${\mathcal{S}}\equiv\{(w,y)\}$ comprise the vector $w$ and outcome $y$. These data are valuable insofar as they make the farmer’s beliefs about $\theta$ more precise. We measure the value of $\mathcal{S}$} via the mean difference
between the prior and posterior variances of $\theta_{1}$ and $\theta_{2}$.
The farmer chooses the vector $w$ that maximizes $\pi({\mathcal{S}})$ subject to the constraint $\lVert w\rVert=1$. Intuitively, he chooses the combination of fertilizers that teaches him as much as possible about their effects. That he can only choose one combination reflects the scarcity and cost of relevant data: our Ugandan setting is one of many where humans must learn from limited data.
@@
The farmer knows the two fertilizers supply equal amounts of nitrogen, a nutrient that helps crops grow. He cannot see or touch nitrogen; it is a mental construct. But he can use his conceptual knowledge of nitrogen to express each fertilizer’s effect $\theta_{k}$ as the sum of a common “nitrogen effect” and an idiosyncratic effect. He encodes these effects by the scalars
@@
These vectors form an orthonormal basis for the Euclidean space $\mathbb{R}^{2}$ containing $\theta$. The common and idiosyncratic effects $\gamma_{1}$ and $\gamma_{2}$ are the coordinates of $\theta$ over this basis.
The farmer knows $v_{1}$ and $v_{2}$, but does not know $\gamma_{1}$ or $\gamma_{2}$. Knowing $v_{1}$ and $v_{2}$ makes learning $\theta$ equivalent to learning $\gamma\equiv(\gamma_{1},\gamma_{2})$. Moreover, since $v_{1}$ and $v_{2}$ are orthonormal, his data $\mathcal{S}$} have value [^4] We derive this expression for $\pi({\mathcal{S}})$ in Section 4.1.
$\pi({\mathcal{S}})=\frac{1}{2}\sum_{k=1}^{2}\mathopen{}\mathclose{{\left(\operatorname{\mathbb{V}}(\gamma_{k})-\operatorname{\mathbb{V}}(\gamma_{k}\mid{\mathcal{S}})}}\right).$
Thus, knowing about nitrogen allows the farmer to reframe his problem from learning about the effect vector $\theta$ to learning about the coordinate vector $\gamma$. He does not have to learn each fertilizer’s effect separately; instead, he can learn the nitrogen and idiosyncratic effects, and extrapolate the overall effects. This is important because he only has one observation $(w,y)$ from which to infer two unknowns $\theta_{1}$ and $\theta_{2}$. Knowing how these unknowns relate (via $v_{1}$ and $v_{2}$) allows him to learn about both at the same time by choosing $w$ appropriately.
The farmer’s choice of $w$ depends on the relative contributions of $\gamma_{1}$ and $\gamma_{2}$ to the prior variances of $\theta_{1}$ and $\theta_{2}$. He knows $\gamma_{1}$ contributes more: the fertilizers’ effects are mostly determined by how much nitrogen they supply. So he assumes $\gamma_{1}$ and $\gamma_{2}$ are independently distributed with variances $\lambda_{1}=\sigma^{2}(1+\rho)$ and $\lambda_{2}=\sigma^{2}(1-\rho)$. The sum
@@
of this sum contributed by $\gamma_{1}$. This share equals $1/2$ when $\rho=0$, in which case $\gamma_{1}$ and $\gamma_{2}$ contribute equally. It equals one in the limit as $\rho\to 1$, in which case only $\gamma_{1}$ contributes. The larger is $\rho$, the more likely is $\theta$ to be close to the one-dimensional subspace of $\mathbb{R}^{2}$ spanned by $v_{1}$.
@@
The value $\pi({\mathcal{S}})$ of the farmer’s data is largest when $w=\pm v_{1}$ and smallest $w=\pm v_{2}$ (see Proposition A2). For example, choosing $w=v_{1}$ makes $y=\gamma_{1}+u$ a “pure signal” of $\gamma_{1}$. This makes $\mathcal{S}$} maximally valuable because it provides information about the component of $\theta$ with the most prior variance, leading to the largest difference between prior and posterior variances. In contrast, choosing $w=v_{2}$ makes $y=\gamma_{2}+u$ a pure signal of $\gamma_{2}$. This makes $\mathcal{S}$} minimally valuable because it provides information about the component of $\theta$ with the least prior variance, leading to the smallest difference between prior and posterior variances. [^5] In general, the data $\mathcal{S}$} are most valuable when they contain information about the components of $\theta$ with the most prior variance. We formalize and prove this claim in Sections 4.14.3.
@@
The farmer attains $\pi^{*}$ by choosing $w=\pm v_{1}$. This choice relies on his conceptual knowledge: he must know $\theta_{1}$ and $\theta_{2}$ share a common nitrogen effect. Without this knowledge, the farmer would have no way to represent $\theta_{k}$ as the sum of components with differential contributions to its prior variance. So he would assume equal contributions (i.e., $\rho=0$) and his data would have maximal value
@@
between $\pi^{*}$ and $\pi^{(0)}$ captures the value of the farmer’s conceptual knowledge: the value of knowing about nitrogen and using this knowledge to make his data more valuable.
The value $\Pi$ of the farmer’s conceptual knowledge is larger when $\rho$ is larger. [^6] We have $\pi^{*}=\frac{(1+\rho)^{2}\sigma^{4}}{2\mathopen{}\mathclose{{\left((1+\rho)\sigma^{2}+\sigma_{u}^{2}}}\right)}\quad\text{and}\quad\pi^{(0)}=\frac{\sigma^{4}}{2(\sigma^{2}+\sigma_{u}^{2})}$ by Proposition A2 and the definition of $\pi^{(0)}$. So $\partial\pi^{*}/\partial\rho>0$ and $\partial\pi^{(0)}/\partial\rho=0$, from which it follows that $\partial\Pi/\partial\rho>0$. , [^7] For example, the correlation $\rho$ will be close to one when the fertilizers supply nitrogen only, and close to zero when their nutrient profiles are very different. Intuitively, if most of fertilizers’ effects come from supplying nitrogen, then the farmer can refine his prior a lot by isolating the nitrogen effect when he tries fertilizers. We formalize this intuition in Section 5, and generalize it to a setting in which $\mathcal{S}$} has arbitrary size and $\theta$ has arbitrary length. In this setting, conceptual knowledge is more valuable when the eigenvalues of the prior variance matrix $\operatorname{\mathbb{V}}(\theta)$ are more spread out (see Theorem 1). This is why raising $\rho$ raises $\Pi$: it raises $\lambda_{1}=(1+\rho)\sigma^{2}$ and lowers $\lambda_{2}=(1-\rho)\sigma^{2}$ without changing their mean $(\lambda_{1}+\lambda_{2})/2=\sigma^{2}$.